golfer hits a 46 g golfball with an initial velocity of 54 m/s at 25°. What are the horizontal and vertical components of momentum imparted to the ball?
1
Expert's answer
2020-04-06T08:55:14-0400

The momentum is


"\vec{p}=m\vec{v}."

The horizontal component:

"p_x=mv_x=mv\text{ cos}\alpha=\\=46\cdot10^{-3}\cdot54\text{ cos}25^\circ=2.3\text{ kg}\cdot\text{m/s}."

The vertical component:


"p_y=mv_y=mv\text{ sin}\alpha=\\=46\cdot10^{-3}\cdot54\text{ sin}25^\circ=2.3\text{ kg}\cdot\text{m/s}."